Skip to main contentSkip to solution

The sum and the product of the mean and variance of binomial distribution are 1.8 and 0.8 respectively. The probability of atleast one success is:

Solution

✅ Correct Option: 3

With mean m=npm=np and variance v=npqv=npq, m+v=1.8m+v=1.8 and mv=0.8mv=0.8, so mm and vv are the roots of t2−1.8t+0.8=0t^2-1.8t+0.8=0, namely 11 and 0.80.8.

The mean exceeds the variance, so np=1np=1, q=0.8q=0.8, p=0.2p=0.2 and n=5n=5.

Hence P(X≥1)=1−(45)5P(X\ge1)=1-\left(\frac{4}{5}\right)^5.

Keyboard Shortcuts

  • Left arrow: Previous question
  • Right arrow: Next question
  • S key: Jump to solution
  • Q key: Jump to question